For a step index fibre with a core of refractive index of 1.54 and cladding refractive index 1.5 calculate the following when fibre is kept in water

b)Numerical aperture

c)Acceptance angle

d)Numerical Aperture in air


1
Expert's answer
2021-05-10T12:17:26-0400

Core refracting index, "\mu_1=1.54"

Cladding refracting index, "\mu_2=1.50"

Refractive index of air, "\mu_o=1"


(b) Numerical Aperture, "(NA)=\sqrt{\mu_1^2-\mu_2^2}"

"=\sqrt{(1.54)^2-(1.50)^2}=0.348"


(c) Acceptance angle, "\theta_a=sin^{-1}(NA)=sin^{-1}(0.348)=20.40\degree"


(d) Numerical aperture in air "=\sqrt{\mu_1^2-\mu_o^2}"

"=\sqrt{(1.54)^2-1^2}=1.17"